In the right triangle shown, mZQ= 60 and QR = 2/3. 60 R- 2v3 How long is PQ? Choose 1 answer:
Answer #1

Cos(60)=1/2=(2*root3)/x. This is because of the trigonometric ratio value of 60 degree.

X=4*(3)^0.5 or 4 root3